Find the empirical formula of a compound that is 48.38 % carbon, 8.12 % hydrogen, and 43.5 % oxygen


1
Expert's answer
2021-01-11T03:50:11-0500

nC = 48.38g C x "\frac{1molC}{12gC}" = 4.03 ÷ 2.71 = 1.5molC

nH = 8.12g H x "\frac{1molH}{1gH}" = 8.12 ÷ 2.71 = 3molH

nO = 43.5g O x "\frac{1molO}{16gO}" = 2.71 ÷ 2.71 = 1molO

now we got a number 1.5 so we multiply by 2 to all

nC = 1.5 ( 2 ) = 3molC

nH = 3 ( 2 ) = 6molH

nO = 1 ( 2 ) = 2molO

now we got empirical formula is

C2H6O2